The sum of two integers is 10, and their product is 24. Find the two integers.
Answer by checkley77(12844)

x+y=10 or x=10-y
xy=24
(10-y)y=24
10y-y^2=24
-y^2+10y-24=0
y^2-10y+24=0
(y-6)(y-4)=0
y-=0
y=6 ans.
x=10-6=4 ans.
y-4=0
y=4 ans.
x=10-4=6 ans.
Proof:
4*6=24